1. Use the data table inside the front cover of Callister or another source of atomic weights and densities to answer this question. A silver wire is 0.4 mm in diameter and 12 cm long. How many atoms it contain?
Answer #1

The wire is cylinderical in shape .

The volume of wire is given as \pi r^{2}h

So volume of the wire will be = 3.14 x (0.2)2 x 120

= 15.086 cubic mm or 1.5086 x 10-2 cubic cm

Density of silver is 10.49 g/cm3

Density = mass / volume

Mass = 10.49 x 1.5086 x 10-2

Mass = 0.158 g

Molar mass of silver = 107.9 g / mol

So 0.158 g = 0.158 / 107.9 = 0.00146 mole

1 mole of Ag has 6.022 x 1023 atoms

So number of atoms present in 0.00146 mole will be 0.00146 x 6.022 x 1023

= 8.79 x 1020 atoms